OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5)

OpenStudy (anonymous):

By slope formula m = (y - y1)/(x - x1), let: (x,y) = (4,6) (x1,y1) = (1,5) So: m = (6 - 5)/(4 - 1) = 1/3 Then, by point-slope formula y - y1 = m(x - x1), letting m = 1/3 and (x1,y1) = (1,5): y - 5 = 1/3 * (x - 1) Since standard form consists of integers, multiply the expression by 3: 3(y - 5) = 1/3 * (x - 1) * 3 3y - 15 = x - 1 So in Ax + By = C form: x - 3y - 1 = -15 x - 3y = -14 [Note that x must be positive integer!]